OpenStudy (anonymous):

Just wondering if I am correct on the format I am doing for such question. Find x (DRAWING WILL BE INSERTED) Please wait about a minute

OpenStudy (anonymous):



OpenStudy (anonymous):

I am thinking that I would have to do half of 41 - which is 20.5 - to get my answer.. Am I correct?

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

Anyone?

OpenStudy (phi):

yes, the "big" (outer triangle) is similar to the smaller top triangle. in this case the smaller triangle's sides are exactly 1/2 the size of the outer, big triangle. that means x is 1/2 of the corresponding side, so 1/2 of 41

OpenStudy (anonymous):

Oh ok! Thank you! And how about one that would have a 3x instead of an x and a 54 instead of a 41?

OpenStudy (anonymous):

Would I make them equal to each other, with the 54 being divided by two?

OpenStudy (anonymous):

\[3x=1/2 (54)\] Would this be correct?

OpenStudy (anonymous):

It still asking me to find x

OpenStudy (phi):

yes, if its the same kind of triangles. 3x = 27 divide both sides by 3 3/3 x = 27/3 x = 9

OpenStudy (anonymous):

Ok, Thank you so much!!!